OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. (2 points) u = <2, -4>, v = <3, -8> 3.0° 6.0° -7.0° 16.0°

OpenStudy (anonymous):

i got -7.0 not sure lol if i got it correct

OpenStudy (nincompoop):

how do you test if it is correct?

OpenStudy (anonymous):

not sure how! actually

OpenStudy (amistre64):

what approach did you take?

OpenStudy (amistre64):

if using vectors, id use the dot product identity

OpenStudy (anonymous):

the last thing i got was (19sqrt365)/ 365 but the decimal form doesnt match my answer

OpenStudy (amistre64):

you want to take a cosine inverse if you did the rest right

OpenStudy (amistre64):

|u| |v| cos(a) = u.v cos(a) = u.v/(|u| |v|) a = inverseCos(u.v/(|u| |v|)) which will either give radians or degrees depending on your calculator setting

OpenStudy (anonymous):

a·b = ax·bx + ay·by = 2 · 3 + (-4) · (-8) = 6 + 32 = 38

OpenStudy (owlcoffee):

You can find the angle between two vectors (also used to find the angle between two lines in an intersecting point) with the equation: \[Cos \alpha =\frac{ x_u x_ v + y_u y_v }{ \sqrt{x_u ^2 + y_u ^2 }\sqrt{x_v ^2 + y_v ^2} }\] of if you prefer: \[Cos \alpha = \frac{ U.V }{ \left| \left| U \right| \right|\left| \left| V \right| \right| }\]

OpenStudy (anonymous):

idk why its so far apart but i used the dot approach

OpenStudy (amistre64):

well, define your lengths and your dot product lets start there and correct if needed

OpenStudy (anonymous):

okay so im confused on the set up then

OpenStudy (amistre64):

@Owlcoffee /cdot makes a nice latex dot \(u\cdot v\)

OpenStudy (owlcoffee):

I'll have that in mind, thanks Amistre.

OpenStudy (amistre64):

u = <2, -4> sqrt(4+16) v = <3, -8> sqrt(9+64) u = <2, -4> v = <3, -8> 6+32 these are our parts right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then a = cos^(-1)(38/(sqrt(20*73)))

OpenStudy (amistre64):

if your calcultor is in radian mode, then it doesn give degrees

OpenStudy (anonymous):

ohh okay let me calculate this

OpenStudy (anonymous):

a=cos^-1 (19/sqrt365)?

OpenStudy (amistre64):

no need to simplify anything, the calculator is capable of working with the stated values

OpenStudy (amistre64):

if your inputting them is awkward, then calculate 20*73 = k take the sqrt of it, write it down if need be or stick it in the calcs memory 38 / k = m now take the inverse cos of m what type of calcuator you using?

OpenStudy (anonymous):

online calculator!

OpenStudy (amistre64):

online, then just do this http://www.wolframalpha.com/input/?i=cos%5E%28-1%29%2838%2F%28sqrt%2820*73%29%29%29

OpenStudy (anonymous):

idk if im inputting the correct thing

OpenStudy (anonymous):

ohh so it would be 6.009!

OpenStudy (amistre64):

yes, or in this case 6.0 seems rnded

OpenStudy (anonymous):

oh i see!! I was forgetting about the radians conversion

OpenStudy (anonymous):

thank you :)!!!!!!!

OpenStudy (amistre64):

your welcome

OpenStudy (anonymous):

would you mind helping me with another! @amistre64